OpenStudy (anonymous):

Tony and Belinda have a combined age of 56. Belinda is 8 more than twice Tony’s age. How old is each?

OpenStudy (anonymous):

tony is 16, belinda is 40

hartnn (hartnn):

let tony's age be x. so Belindas age = 56-x. Belinda is 8 more than twice Tony’s age. so 56-x = 2x+8 can u solve this and find x ?

OpenStudy (farmdawgnation):

@Turner Please provide more than just the answer. Try to work with the asker as hartnn has.

OpenStudy (anonymous):

56 - 8 = 48 48 / 3 = 16 (tony's age) 16 x 2 + 8 = 40 (belinda's age)

OpenStudy (anonymous):

thanks :p

OpenStudy (anonymous):

np (: